OpenStudy (anonymous):

Find the area of three-fifths of a circle with radius 15 ft.

OpenStudy (jmark):

Given Area of three-fifths of a circle with radius 15ft Area of circle with three-fifths of a circle = 3/5*π*r2 = (3/5) *3.14*(15)2 = 423.9 = 424 ft2 CBSE Syllabus for Class 10 2013-2014 http://goo.gl/zX1Xd8